OpenStudy (anonymous):

a⁵ x a x a⁰ x a³ A.) a³ B.) a² C.) a¹⁵ D.) 1

OpenStudy (uri):

Add all the powers.

OpenStudy (uri):

I can't really see what's the power of the last one.. @benetsaidso

OpenStudy (unklerhaukus):

\[x^n=\overbrace{x\times x\times x\times\dots\times x}^{n \text{ times}}\]

OpenStudy (anonymous):

The last one is -3

OpenStudy (unklerhaukus):

\[x^n\times x^m=\overbrace{x\times x\times x\times\dots\times x}^{n \text{ times}}\times\overbrace{x\times x\times x\times\dots\times x}^{m \text{ times}}\\\qquad\quad=\overbrace{x\times x\times x\times\dots\times x}^{n+m \text{ times}}\\\qquad\quad=x^{n+m}\]

OpenStudy (uri):

@benetsaidso What do you think the answer will be?

OpenStudy (anonymous):

I really don't know.. That's why I posted it.

OpenStudy (uri):

I just told you to add all the powers..and ofcourse if it's -3 you will subract it.

OpenStudy (unklerhaukus):

\[x^n\times x^m\times x^p\times x^{-q}=x^{n+m+p-q}\]

OpenStudy (anonymous):

B ?

OpenStudy (akashdeepdeb):

Okay so when you add all the powers you get a^5+1+0-3 Thus a^3 Thus the answer is A) a^3! :D

OpenStudy (anonymous):

A ? :S

OpenStudy (uri):

a^5=5 power,a= a^1=1 power a^0 =0 and a^-3=-3 power 5+1+0-3 6-3 =3

OpenStudy (unklerhaukus):

note that \[x^1=\overbrace{x}^{\text{1 time}}\] \[x^0=\overbrace{1}^{\text{0 times}}\]

OpenStudy (uri):

a^3